OpenStudy (anonymous):

The second term in a geometric sequence is 20. The fourth term in the same sequence is 45/4, or 11.25. what is the common ratio in this sequence?

OpenStudy (campbell_st):

well in a geometric sequence a term can be found by using \[a_{n} = ar^{n -1}\] so you know \[20 = ar\] and you also know \[11.25 = ar^3\] the 4th term can be rewritten as \[11.25 = (ar) \times r^2\] which means its written in terms of the 2nd term so \[11.25 = 20r^2\] just solve the equation for r

OpenStudy (anonymous):

thanks